PLEASE HELP AND SHOW WORK, please highlight answers, THANK YOU!!

13.Direct Labor Variances Bellingham Company produces a product that requires 3 standard direct labor hours per unit at a standa

14.Net Present Value A project has estimated annual net cash flows of $5,000 for two years and is estimated to cost $30,000. Ass

15.Internal Rate of Return A project is estimated to cost $159,648 and provide annual net cash flows of $48,000 for six years. P

16.Zoe Corporation has the following information for the month of March: Cost of direct materials used in production $15,139 Dir

17.A quality control activity analysis indicated the following four activity costs of a hotel: Inspecting cleanliness of rooms $
Answer #1

Direct labor rate variance = (SR-AR)AH = (10-10.40)*11500 = 4600 U

Direct labor time variance = (SH-AH)SR = (3900*3-11500)*10 = -2000 F

Direct labor time variance = 2600 U

Note : Pls post each que individually as HOMEWORKLIB guidelines